OpenStudy (anonymous):

Let E be the error whe the taylor polynomial T(x) = x - (x^3/3!) centered about x=0, is used to approximate f(x) = sinx at x=0.5. Which is true?

OpenStudy (anonymous):

\[a) 0.0000 > \left| E \right|\]\[b) 0.0001 < \left| E \right| < 0.0003\] c) \[c) 0.0003 < \left| E \right| < 0.0005\]

OpenStudy (anonymous):

Please show work b/c thats wats important not really the asnwer

OpenStudy (anonymous):

cuz i know the answer, i just dont know how to show work